# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ Abstract
We study covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the Generalized Wendland ($\mathcal{GW}$) and Matérn ($\mathcal{MT}$) models. We derive sharp validity conditions, providing a complete characterization of the admissible parameter space, and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics.
To resolve this issue, we introduce a parsimonious compactly supported subclass selected via a maximum integral range criterion. The resulting hypergeometric model can be viewed as a structural refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the $\mathcal{MT}$ model as a limit case.
We further establish strong consistency and asymptotic normality of the maximum likelihood estimator of the associated microergodic parameter under fixed-domain asymptotics. Simulation experiments and a real-data application to climate data illustrate the finite-sample behavior and practical performance of the proposed model.
@@ 1 Introduction
For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework. Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths. A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006). Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
@@
We focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance model introduced by Emery and Alegría (2022). This highly flexible class includes both $\mathcal{MT}$ and $\mathcal{GW}$ as special or asymptotic cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024). While sufficient validity conditions were derived in Emery and Alegría (2022), a complete characterization of the parameter space and a detailed inferential analysis remain open.
@@
This paper provides several theoretical and methodological contributions. First, we refine the validity conditions of the $\mathcal{GH}$ model. We show that the sufficient conditions of Emery and Alegría (2022) are also necessary in the parameter range where the two sets of conditions agree. Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense. The parameter-to-correlation mapping is not injective, and this structural non-identifiability must be distinguished from the well-known inferential limitations under fixed-domain asymptotics, where Gaussian measures may remain equivalent on bounded domains. In such settings, only microergodic parameter combinations are consistently estimable.
@@
To resolve the structural non-identifiability, we introduce an optimality-driven parsimonious subclass of $\mathcal{GH}$. For fixed smoothness and compact support, we select the member that maximizes the isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$. Since the integral range coincides with the spectral density at zero frequency, this criterion selects the admissible model with maximal low-frequency spectral power, i.e., the strongest large-scale dependence compatible with the imposed constraints. The integral range provides a global and threshold-free measure of dependence (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function. The proposed selection principle is intrinsic to the model class and does not depend on sampling design or loss functions.
@@
The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper. It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility. Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension. In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
@@
A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model in terms of model fit, prediction accuracy, and computational efficiency for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
@@ 2 Background: the Matérn and Generalized Wendland correlation models
We denote $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D$ of $\mathbb{R}^{d}$, with stationary isotropic correlation function $C:\mathbb{R}^{d}\to\mathbb{R}$. We consider the class $\Phi_{d}$ of continuous mappings $\phi:[0,\infty)\to\mathbb{R}$ with $\sigma^{2}:=\phi(0)=1$, such that
${\rm corr}\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),$
with $\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s}$ and $\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$, and $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
@@
    ${\cal MT}_{\nu,\alpha}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{\alpha}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{\alpha}\right),\qquad x>0,$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{\alpha,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,\alpha}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{\alpha^{d}}{(1+\alpha^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein (1999) show that ${\cal MT}_{\nu,\alpha}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $\alpha$ are strictly positive.
    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF. In particular, for any integer $l=0,1,\ldots$, the sample paths are $l$-times mean-square differentiable (in any direction) if and only if $\nu>l$. In the special case $\nu=l+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $l$.
@@
    $\begin{split}{\cal GW}_{\kappa,\mu,\beta}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{\beta^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{\beta^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $\beta>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,\beta}(z)=L\beta^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(z\beta)^{2}}{4}\Big),\quad z>0,$ (5)
    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\nu+1)\Gamma(2\nu+d)\Gamma(\nu)}{\Gamma\left(\nu+{d}/{2}\right)\Gamma(\mu+2\nu+d+1)\Gamma(2\nu)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
@@
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
@@
      $\mathcal{GW}_{\kappa,\mu,\beta}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
@@
    The ${\cal GW}_{\kappa,\mu,\beta}$ model, as in the ${\cal MT}_{\nu,\alpha}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $l=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,\beta}$ model are $l$ times differentiable, in any direction, if and only if $\kappa>l-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
@@
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(\beta,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,\beta}(x),\quad\kappa\geq-0.5,$ (6)
    where $g(\beta,\mu,\kappa)=\beta\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
@@ 3 The Gauss hypergeometric model and the identifiability problem
  1. (a)
    $\delta>\frac{d}{2}$;
  2. (b)
    $2(\beta-\delta)(\gamma-\delta)\geq\delta$;
  3. (c)
    $\beta+\gamma\geq 3\delta+\frac{1}{2}$;
@@
Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases. For instance,
@@
Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense. Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable—the parameter-to-correlation map is not injective. In particular,
$\mathcal{GH}_{\delta,\beta,\gamma,a}=\mathcal{GH}_{\delta,\gamma,\beta,a},$
@@
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$. If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, if and only if
@@
Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases. However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged. This issue can be observed under both increasing and fixed domain scenario. To this end, in the next Section we assume that the second shape parameter ($\gamma$) is a function of the first ($\beta$).

### v3

@@ Abstract
We study isotropic covariance functions in the Gauss hypergeometric ($\mathcal{GH}$) class, a flexible family that encompasses the celebrated Matérn ($\mathcal{MT}$) and Generalized Wendland ($\mathcal{GW}$) models. We refine validity conditions for positive definiteness in the $\mathcal{GH}$ family and clarify two distinct sources of non-identifiability: a structural one, due to an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, where different parameter values may induce equivalent Gaussian measures and only microergodic combinations are consistently estimable.
To remove the structural ambiguity, we first restrict the parameter space to obtain a unique representation of the correlation function. We then introduce a parsimonious compactly supported subclass selected by a maximum integral range criterion, leading to a new hypergeometric model, denoted $\mathcal{H}$, with simple validity conditions and sparse covariance matrices. The proposed model can be viewed as a canonical refinement of the $\mathcal{GW}$ family and admits compact-support reparameterizations that recover the Matérn model as a limit case.
Under fixed-domain asymptotics, we establish strong consistency and asymptotic normality for likelihood-based estimation of the associated microergodic parameter, including cases where nuisance covariance parameters are estimated over compact sets. Simulation experiments and a climate-data application illustrate finite-sample behavior and show that $\mathcal{H}$ can improve upon $\mathcal{MT}$ while achieving performance comparable to $\mathcal{GW}$ in terms of fit, prediction, and computational efficiency for both Gaussian and Tukey-hh random fields.
@@ 1 Introduction
For RFs on Euclidean spaces, isotropic covariance (kernel) functions— depending only on Euclidean distance—provide a tractable and widely used modeling framework. Among parametric isotropic models, the Matérn ($\mathcal{MT}$) class is arguably the most prominent (Porcu et al., 2024), owing to its smoothness parameter, which continuously controls the mean-square differentiability of the sample paths.
A compactly supported alternative is the generalized Wendland ($\mathcal{GW}$) family (Gneiting, 2002; Bevilacqua et al., 2019), which also allows for continuous control of smoothness while yielding sparse covariance matrices and substantial computational gains (Duff et al., 1989; Davis, 2006) when working with covariance matrices. Connections between $\mathcal{MT}$ and $\mathcal{GW}$ models have been extensively studied (Bevilacqua et al., 2019, 2022).
@@
In this paper, we focus on the Gauss hypergeometric ($\mathcal{GH}$) covariance function introduced by Emery and Alegría (2022). This highly flexible class includes the Matérn ($\mathcal{MT}$) and generalized Wendland ($\mathcal{GW}$) models as special cases and has been recognized as a prominent extension of the Matérn family (Porcu et al., 2024). Although Emery and Alegría (2022) derived convenient sufficient conditions for positive definiteness, a sharper description of the admissible parameter space and a systematic inferential analysis are still lacking. The aim of this paper is to address these issues.
@@
This paper provides several theoretical and methodological contributions. First, we sharpen the validity conditions of the $\mathcal{GH}$ class by deriving explicit constraints that characterize positive definiteness over a substantial portion of the parameter space, thereby improving the results of Emery and Alegría (2022). Second, we clarify that the $\mathcal{GH}$ family is not identifiable in the classical parametric sense: the parameter-to-correlation mapping is not injective due to an intrinsic symmetry in the shape parameters. We distinguish this structural non-identifiability from the well-known inferential limitations under fixed-domain asymptotics, where different parameter values may induce equivalent Gaussian measures and only microergodic parameter combinations are consistently estimable.
To address the structural non-identifiability, we first impose a restriction on the parameter space of the $\mathcal{GH}$ model so that each admissible parameter set corresponds to a unique correlation function.
@@
Once this structural issue is resolved, we pursue a separate goal: identifying a parsimonious and canonical compactly supported subclass within $\mathcal{GH}$. To this end, for fixed smoothness and compact support we adopt an integral-range criterion, selecting the subclass that can attain the largest isotropic integral range, that is, the total mass of the correlation over $\mathbb{R}^{d}$. Since the integral range is proportional to the spectral density at zero frequency, this criterion favors models with the largest low-frequency spectral power, namely the strongest large-scale dependence compatible with the imposed constraints. The integral range thus provides a global and threshold-free measure of dependence. (Vanmarcke, 1983; Lantuéjoul, 2002; Hristopulos and Zukovic, 2011); in dimension $d=1$, it reduces to the area under the correlation function.
@@
The resulting hypergeometric ($\mathcal{H}$) model constitutes the main contribution of this paper. It provides a canonical and parsimonious representative of the $\mathcal{GH}$ class, preserving compact support, continuous smoothness parameterization, and modeling flexibility and can be viewed as a refinement of the ($\mathcal{GW}$) family. Notably, the $\mathcal{H}$ model satisfies simple and explicit necessary and sufficient validity conditions, reducing to a single lower bound on the shape parameter that is independent of the spatial dimension. In addition, it admits a beta-type scale-mixture representation and may be interpreted as a generalization of the Euclid’s hat model (Gneiting, 1999), encompassing the classical spherical covariance functions widely used in geostatistics (Chilès and Delfiner, 2012).
@@
Finally, for a fixed and known smoothness parameter, we establish strong consistency and asymptotic normality for likelihood-based estimation of the microergodic parameter under the proposed $\mathcal{H}$ model, including settings where scale and shape covariance parameters are estimated over compact sets. These results extend, in the hypergeometric framework, the fixed-domain asymptotic theory developed for compactly supported models in Bevilacqua et al. (2019).
A simulation study investigates finite-sample behavior under increasing- and fixed-domain scenarios, and a real-data analysis of climate data illustrates that the proposed $\mathcal{H}$ model can outperform the $\mathcal{MT}$ model and achieve performance comparable to the $\mathcal{GW}$ model, in terms of model fit, prediction accuracy, and computational efficiency, for both Gaussian and Tukey-$h$ RFs (Xu and Genton, 2017).
@@ 2 Background: the Matérn and Generalized Wendland correlation models
We denote by $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ a zero-mean Gaussian RF on a bounded set $D\subset\mathbb{R}^{d}$ with stationary isotropic covariance function
${\rm Cov}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=\sigma^{2}\,C(\boldsymbol{h})=\sigma^{2}\,\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{h}=\boldsymbol{s}^{\prime}-\boldsymbol{s},$
where $\sigma^{2}>0$ is the variance parameter and $\phi:[0,\infty)\to\mathbb{R}$ is a correlation function satisfying $\phi(0)=1$. We consider the class $\Phi_{d}$ of continuous mappings $\phi$ such that $\phi(\|\cdot\|)$ is positive semidefinite on $\mathbb{R}^{d}$ (equivalently, $\phi\in\Phi_{d}$). In particular,
${\rm corr}\!\left(Z(\boldsymbol{s}),Z(\boldsymbol{s}^{\prime})\right)=C(\boldsymbol{h})=\phi(\|\boldsymbol{h}\|),\qquad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D,$
with $\|\cdot\|$ denoting the Euclidean norm. Gaussian RFs with such covariance functions are called weakly stationary and isotropic.
@@
    ${\cal MT}_{\nu,a}(x)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{x}{a}\right)^{\nu}{\cal K}_{\nu}\left(\frac{x}{a}\right),\qquad x>0,$ (2)
    where ${\cal K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz and Stegun, 1970, 9.6.22). The associated spectral density, $\widehat{{\cal MT}}_{a,\nu}$, is given by
    $\widehat{{\cal MT}}_{\nu,a}(z)=\frac{\Gamma(\nu+d/2)}{\pi^{d/2}\Gamma(\nu)}\frac{a^{d}}{(1+a^{2}z^{2})^{\nu+d/2}},\qquad z\geq 0.$ (3)
    Arguments in Stein (1999) show that ${\cal MT}_{\nu,a}$ belongs to $\Phi_{\infty}$ when both the smoothness parameter $\nu$ and the scale parameter $a$ are strictly positive.
    A key feature of the Matérn model is the smoothness parameter $\nu$, which determines the mean-square differentiability of the sample paths of the associated Gaussian RF. In particular, for any integer $m=0,1,\ldots$, the sample paths are $m$-times mean-square differentiable (in any direction) if and only if $\nu>m$. In the special case $\nu=m+\tfrac{1}{2}$, the correlation function reduces to an exponential term multiplied by a polynomial of degree $m$.
@@
    $\begin{split}{\cal GW}_{\kappa,\mu,a}(x)&=\frac{\Gamma(\kappa)\Gamma(2\kappa+\mu+1)}{\Gamma(2\kappa)\Gamma(\kappa+\mu+1)2^{\mu+1}}\left(1-\frac{x^{2}}{a^{2}}\right)^{\kappa+\mu}_{+}\\ &\times{}_{2}F_{1}\left(\frac{\mu}{2},\frac{\mu+1}{2};\kappa+\mu+1;\left(1-\frac{x^{2}}{a^{2}}\right)_{+}\right),\quad x\geq 0,\end{split}$ (4)
    with $a>0$ a compact support parameter, $\kappa>-\frac{1}{2}$ a smoothness parameter and $\mu$ a shape parameter, and where $(\cdot)_{+}$ stands for the positive part function and $_{2}F_{1}$ for the Gauss hypergeometric function (Abramowitz and Stegun, 1970, 15.1.1). The associated spectral density is given by:
    $\widehat{{\cal GW}}_{\kappa,\mu,a}(z)=La^{d}\mathstrut_{1}F_{2}\Big(\lambda;\lambda+\frac{\mu}{2},\lambda+\frac{\mu}{2}+\frac{1}{2};-\frac{(za)^{2}}{4}\Big),\quad z>0,$ (5)
    where $_{1}F_{2}$ is a generalized hypergeometric function, $\lambda=(d+1)/2+\kappa$ and $L=\frac{2^{-d}\pi^{-\frac{d}{2}}\Gamma(\mu+2\kappa+1)\Gamma(2\kappa+d)\Gamma(\kappa)}{\Gamma\left(\kappa+{d}/{2}\right)\Gamma(\mu+2\kappa+d+1)\Gamma(2\kappa)}$ is a normalization constant.
    For a given $\kappa>-\frac{1}{2}$, the validity conditions for $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ follow as a special case of Theorem 2 by setting $l=\frac{1}{2}$:
@@
      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{d}$ if and only if $\mu\geq\kappa+\frac{d+1}{2}$, whenever $\kappa\geq 0$ or $d\geq 2$;
@@
      $\mathcal{GW}_{\kappa,\mu,a}\in\Phi_{1}$ if $\mu\geq\frac{\sqrt{8\kappa+9}-1}{2}$, and only if $\mu\geq\kappa+1$, for $\frac{1}{2}<\kappa<0$.
@@
    The ${\cal GW}_{\kappa,\mu,a}$ model, as in the ${\cal MT}_{\nu,a}$ case, allows parameterizing in a continuous fashion the mean squared differentiability of the underlying Gaussian RF and its associated sample paths through the smoothness parameter $\kappa$. Specifically, for any integer $m=1,2\ldots$, the sample paths of the ${\cal GW}_{\kappa,\mu,a}$ model are $m$ times differentiable, in any direction, if and only if $\kappa>m-0.5$ and for $0\leq\kappa\leq 0.5$ they are not differentiable.
@@
    $\lim_{\mu\to\infty}{\cal GW}_{\kappa,\mu,g(a,\mu,\kappa)}(x)={\cal MT}_{\kappa+1/2,a}(x),\quad\kappa>-0.5,$ (6)
    where $g(a,\mu,\kappa)=a\left(\frac{\Gamma(\mu+2\kappa+1)}{\Gamma(\mu)}\right)^{\frac{1}{1+2\kappa}}$, with uniform convergence on $0,\infty)$.
@@ 3 The Gauss hypergeometric model and the identifiability problem
$(a)\ \delta>\frac{d}{2},\qquad(b)\ 2(\beta-\delta)(\gamma-\delta)\geq\delta,\qquad(c)\ \beta+\gamma\geq 3\delta+\frac{1}{2}\,$,
@@
Furthermore, it is not difficult to see that the $\cal G$H} model includes the $\cal GW$} and the $\cal MT$} models as special cases (Emery and Alegría, 2022). For instance:
@@
Although the $\cal G$H} family is highly flexible, it is not identifiable in the usual parametric sense. Indeed, even under increasing-domain asymptotics—where one would in principle expect all parameters to be estimable (Mardia and Marshall, 1984)—the parameter-to-correlation map is not injective. In particular,
$\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x),$
@@
Let $P(\sigma_{i}^{2}{\cal GH}_{\delta,\beta_{i},\gamma_{i},a_{i}})$, $i=0,1$, be two zero-mean Gaussian measures and let $\delta>\frac{d}{2}$. If $\frac{d+1}{2}+3\delta<\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0}),$ and $2(\beta_{i}-\delta)(\gamma_{i}-\delta)\geq\delta.$ for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{1}^{2}{\cal GH}_{\delta,\beta_{1},\gamma_{1},a_{1}})$ and $P(\sigma_{0}^{2}{\cal GH}_{\delta,\beta_{0},\gamma_{0},a_{0}})$ are equivalent on the paths of $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$, if and only if
@@
Summarizing, the $\mathcal{GH}$ correlation model is highly flexible and includes the $\mathcal{GW}$ and $\mathcal{MT}$ models as special cases. However, it is not identifiable in the usual parametric sense because swapping $(\beta,\gamma)$ leaves the correlation unchanged. This issue can be observed under both increasing and fixed domain scenario.